OpenStudy (anonymous):

The hypotenuse of a right triangle is 7 inches, and one of the legs is ( ) inches. Find the length of the other leg.

OpenStudy (anonymous):

( \[\sqrt{13}\])

OpenStudy (anonymous):

Pythagorean Theorem. a^2 + b^2 = c^2. You have two values, plug them in and solve for the third.

OpenStudy (anonymous):

49-13=36 \[\sqrt{36}\]=6

OpenStudy (anonymous):

Please don't just give the answer...